OpenStudy (anonymous):

please help what is the solution to the system of equations y=2x+6 3x-y=6

OpenStudy (anonymous):

@undeadknight26

pooja195 (pooja195):

\[ 3x-(2x+6)=6\]

OpenStudy (anonymous):

thank you so much can you check my other questions

OpenStudy (anonymous):

i want to see it i got them right

pooja195 (pooja195):

This was only the first step xD

OpenStudy (anonymous):

oh what are zee others

OpenStudy (anonymous):

would it be 6,6

pooja195 (pooja195):

hold on im checking the work

OpenStudy (anonymous):

ok

pooja195 (pooja195):

thats not right

OpenStudy (anonymous):

whats not right?

pooja195 (pooja195):

Ok show me what you did type it out

OpenStudy (anonymous):

i don't know how to do it

OpenStudy (anonymous):

s why i asked you

OpenStudy (anonymous):

but ill try

pooja195 (pooja195):

\[3x−2x+6=6~=1x+6\]

OpenStudy (anonymous):

what are the ?'s for

pooja195 (pooja195):

refresh your page

OpenStudy (anonymous):

why do i have to be sooo dumb

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

so x = 0

pooja195 (pooja195):

Yep :D your not dumb :) dont say that

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

so how would i find out y

pooja195 (pooja195):

now \[y=2(0)+6\]

OpenStudy (anonymous):

y = 6

OpenStudy (anonymous):

so its 0,6

pooja195 (pooja195):

yep

OpenStudy (anonymous):

yay thank you sooo much again

pooja195 (pooja195):

no probs friend :D

OpenStudy (anonymous):

hold on ill tell you what i get

OpenStudy (anonymous):

it says the answer was 12,30

pooja195 (pooja195):

what method were u taught to use?

OpenStudy (anonymous):

um substitute

OpenStudy (anonymous):

what is the solution to the system of equations x+y=2 x-y=4

pooja195 (pooja195):

ok so you want to isolate the x from the first equation so subtract y

pooja195 (pooja195):

\[x=y+4\]

pooja195 (pooja195):

i meant add y

pooja195 (pooja195):

@Nnesha help me out